A population N = 5 consists of the following values: 11, 13, 15, 17, and 19. Estimate the population mean (u) using a random variable of size n = 3

1
Expert's answer
2022-05-08T13:50:52-0400

a. We have population values 11, 13, 15, 17, and 19, population size N=4 and sample size n=3.

Mean of population "(\mu)" = 

"\dfrac{11+13+15+17+19}{5}=15"


The number of possible samples which can be drawn without replacement is "\dbinom{5}{3}=10."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 &11,13,15 & 39/3 \\
 \hdashline
 2 & 11,13,17 & 41/3 \\
 \hdashline
 3 & 11,13,19 & 43/3 \\
 \hdashline
 4 & 11,15,17 & 43/3 \\
 \hdashline
 5 & 11,15,19 & 45/3 \\
 \hdashline
 6 & 11,17,19 & 47/3 \\
 \hdashline
 7 & 13,15,17 & 45/3 \\
 \hdashline
 8 & 13,15,19 & 47/3 \\
 \hdashline
 9 & 13,17,19 & 49/3 \\
 \hdashline
 10 & 15,17,19 & 51/3 \\
\end{array}"





"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X})\\ \hline
 39/3 & 1/10 & 39/30\\
 \hdashline
 41/3 & 1/10 & 41/30\\
 \hdashline
 43/3 & 2/10 & 86/30\\
 \hdashline
 45/3 & 2/10 & 90/30\\
 \hdashline
 47/3 & 2/10 & 94/30\\
 \hdashline
 49/3 & 1/10 & 49/30\\
 \hdashline
 51/3 & 1/10 & 51/30\\
 \hdashline

\end{array}"


Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=\dfrac{450}{30}=15=\mu"